a room is half as long again as its broad and the area of its floor is 98 sq m . find its lenght and breadth

Grade:12th pass

1 Answers

Vikas TU
6 years ago
a room is half as long again as its broad:
 
breadth, b= h/2 = l
where h is the length.
h/2*h/2 = 98sqm
h^2 = 98*4
h = 2root98
 
l = b = h/2 = root98.